Chuck, a single taxpayer, earns $86,000 in taxable income and $20,000 in interest from an investment in City of Heflin bonds. (Use the U.S. tax rate schedule.)

Required:

  1. If Chuck earns an additional $53,000 of taxable income, what is his marginal tax rate on this income?
  2. What is his marginal rate if, instead, he had $53,000 of additional deductions?

2019 Tax Rate Schedules

Individuals

Schedule X-Single

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $510,300 $46,628.50 plus 35% of the excess over $204,100
$510,300 $153,798.50 plus 37% of the excess over $510,300
Answer #1
Tax on $86000 14814.5 =14382.5+(86000-84200)*24%
a
Taxable income 139000 =86000+53000
Tax on $139000 27534.5 =14382.5+(139000-84200)*24%
Marginal tax rate 24.00% =(27534.5-14814.5)/53000
b
Taxable income 33000 =86000-53000
Tax on $33000 3766 =970+(33000-9700)*12%
Marginal tax rate 20.85% =(14814.5-3766)/53000